# Finite-Time Analysis of Simultaneous Double Q-learning

### v1

@@ 4.3 Finite-time error bounds
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\rho^{k-4}k^{4}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}.$ (7)
@@
The proof is given in Section C.5 of Appendix. The bound in (7) can be converted to more interpretable form presented below.
@@
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\frac{\rho^{-4}(-8)^{4}}{(\ln(\rho))^{4}}\rho^{\frac{-4}{\ln(\rho)}}\rho^{k/2}.$ (8)
@@
The proof is given in Section C.6 of Appendix. The remaining parts of the paper are devoted to brief sketches of the proofs.
@@ 5.1 Switching system model
$w_{k}^{A}$ $=(e_{a_{k}}\otimes e_{s_{k}})r_{k}+\gamma(e_{a}\otimes e_{s})(e_{s_{k}^{\prime}})^{T}\Pi_{Q_{k}^{B}}Q_{k}^{A}-(e_{a_{k}}\otimes e_{s_{k}})(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}$
@@
$w_{k}^{B}$ $=(e_{a_{k}}\otimes e_{s_{k}})r_{k}+\gamma(e_{a}\otimes e_{s})(e_{s_{k}^{\prime}})^{T}\Pi_{Q_{k}^{A}}Q_{k}^{B}-(e_{a_{k}}\otimes e_{s_{k}})(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}$
@@
Next, using the optimal Bellman equation $(\gamma DP\Pi_{Q^{*}}-D)Q^{*}+DR=0$ with (5.1), one can obtain
@@
which is a linear switching system with an extra affine terms, $\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q^{*}}Q^{*}\}$ and $\alpha D\{\gamma P\Pi_{Q_{k}^{A}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}\}$, and the stochastic noises, $w_{k}^{A}$ and $w_{k}^{B}$  [12].
The main difficulty in analysing the system is due to the extra affine term and the stochastic noise, without which the finite-time analysis is straightforward, as stability of the system matrix can be analyzed directly. However, with the affine term, the analysis becomes more challenging. To address this difficulty, we introduce the lower and upper comparison systems as in [12], which enable easier analysis.
@@ 5.2 Upper comparison system
$Q_{k+1}^{A_{U}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)(Q_{k}^{A_{U}}-Q^{*})+\alpha w_{k}^{A},\quad Q_{0}^{A_{U}}-Q^{*}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$
$Q_{k+1}^{B_{U}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D)(Q_{k}^{B_{U}}-Q^{*})+\alpha w_{k}^{B},\quad Q_{0}^{B_{U}}-Q^{*}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert}.$ (12)
The above systems are switching systems, which have system matrices $I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D$ and $I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D$. These matrices switch according to the changes of $Q_{k}^{A}$ and $Q_{k}^{B}$.
We can prove that the trajectory of the upper comparison system bounds that of the original system from above.
@@ 5.3 Lower comparison system
$Q_{k+1}^{A_{L}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
$\quad+\alpha w_{k}^{A},\quad Q_{0}^{A_{L}}-Q^{*}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$
$Q_{k+1}^{B_{L}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})$
$\quad+\alpha w_{k}^{B},\quad Q_{0}^{B_{L}}-Q^{*}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$ (13)
where the stochastic noises $w_{k}^{A}$ and $w_{k}^{B}$ are identical to the original system (5.1).
As before, we can prove that the trajectory of the lower comparison system bounds that of the original system from below.
@@
The proof is in Section 5.2 of Appendix. The lower comparison system (5.3) can be seen as a linear system with the states $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$ and the system matrix $I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D$. Moreover, it also includes the extra terms, $\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$ and $\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})$, which can be seen as external disturbances. To derive a finite-time error bound, one needs to establish bounds on the error $Q_{k}^{A}-Q_{k}^{B}$ first. Therefore, in the next subsections, we introduce an error system.
@@ 5.4 Error system
$Q_{k+1}^{\text{err}}$ $=(I-\alpha D)Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$ (14)
@@
However, the error system in (14) has the affine term $\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}$ similar to the original double $Q$-learning or $Q$-learning. Therefore, one can imagine that its convergence can be proved using similar techniques as in the $Q$-learning analysis [11, 12, 13]. In particular, one can derive the upper and lower comparison systems of the error system, and then prove the overall convergence using the fact that the lower comparison system is linear. However, for the error system (14), similar ideas cannot be applied because the lower comparison system of the error system is a switching system. For this reason, we will use a different approach.
In particular, we will first consider an upper comparison system of the error system, and then, derive a lower comparison system of the upper comparison system, which is linear. Then, we will consider a lower comparison system of the error system. Because the lower comparison system is a switching system, we will derive a subtraction system which can be obtained by subtracting the error lower comparison system from the error upper comparison system.

### v2

@@ 4.3 Finite-time error bounds
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\mathrm{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\rho^{k-4}k^{4}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}.$ (8)
@@
The proof is given in C.5. The bound in (8) consists of two terms with distinct interpretations. The second term decays exponentially fast as $k$ increases, since $\rho\in(0,1)$, and therefore vanishes exponentially. The first term represents a constant error term that depends on the step size $\alpha$ and the minimum state–action visitation probability $d_{\mathrm{min}}$. By choosing a sufficiently small step size, this term can be made arbitrarily small. Moreover, $d_{\mathrm{min}}$ characterizes the level of exploration in the learning process: under uniform exploration, $d_{\mathrm{min}}$ is large, and it leads to a smaller error bound, whereas non-uniform or poor exploration results in a smaller $d_{\mathrm{min}}$ and thus a larger error bound. The bound in (8) can be converted to more interpretable form presented below.
@@
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\mathrm{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}\frac{\rho^{-4}(-8)^{4}}{(\ln(\rho))^{4}}\rho^{\frac{-4}{\ln(\rho)}}\rho^{k/2}.$ (9)
@@
The proof is given in C.6.
4.3.1 Comparative convergence analysis
We summarize in Table 1 the sample complexities of representative double $Q$-learning and $Q$-learning algorithms, each derived under distinct assumptions and observation models. The comparison is organized along three key dimensions: (i) the sampling model, which distinguishes between i.i.d. and non-i.i.d. data generation; (ii) the coverage condition, which characterizes how sufficiently all state–action pairs are explored; and (iii) the step-size rule, which determines whether the learning rate is constant or diminishing over time. Specifically, the coverage condition takes one of three forms: the cover-time assumption, which requires that every state–action pair be visited at least once within a finite time window; the infinite-time covering assumption, which ensures that every state–action pair is visited infinitely often over time; and assumes sampling from a stationary distribution with stochastic coverage, meaning that each state–action pair has a strictly positive sampling probability. For completeness, asymptotic convergence results are also included to provide a broader perspective on the overall convergence landscape.
Double $Q$-learning. Earlier works such as [18, 19] analyze the convergence properties of double $Q$-learning in a non-i.i.d. setting under cover-time assumptions. Specifically, [18] employs a polynomially decaying step-size. In contrast, [19] adopts a constant step-size. Our finite-time framework, by comparison, accommodates a general step-size $\alpha\in(0,1)$. Moreover, we instead assume an i.i.d. sampling with stochastic coverage.
[…]
Discussion. Overall, the presented methods should be viewed as complementary rather than competing approaches. Each analysis is conducted under distinct assumptions and observation models, and thus emphasizes different aspects of the convergence behavior of $Q$-learning and double $Q$-learning. Our SDQ analysis does not aim to outperform existing analysis of $Q$-learning or double $Q$-learning in a theoretical sense, but rather to provide a unified interpretation based on a switching-system viewpoint and to establish finite-time expected error bounds within that framework. It should be noted that, under identical initialization ($Q_{0}^{A}=Q_{0}^{B}$), the SDQ update exactly reduces to the standard $Q$-learning algorithm, as discussed in Section 4.1. Hence, no theoretical improvement over $Q$-learning can be expected in this case. The contribution of this work lies not in achieving a tighter asymptotic rate, but in offering a generalized and control-theoretically interpretable framework that unifies $Q$-learning and double $Q$-learning within the same dynamical system formulation. Empirically, as presented in Section 4.2, the simultaneous update structure of SDQ tends to yield faster stabilization under random initialization, which supports the practical relevance and theoretical motivation of this study.
Furthermore, Theorem 4.2 primarily focuses on finite-time estimation accuracy rather than bias analysis, the bias-reduction effect of SDQ arises implicitly from its cross-evaluation structure, each estimator uses the other’s greedy action as the target, thereby reducing the correlation between target selection and estimation noise, a mechanism analogous to that of standard double $Q$-learning. Therefore, the proposed analysis should be regarded as a complementary and explanatory framework rather than a competing algorithmic enhancement. The remaining parts of the paper are devoted to brief sketches of the proofs.
Table 1: Comparative analysis of several results: $t_{\rm mix}$ is the mixing time; $t_{\rm cover}$ is the cover time; $w\in(0,1)$ is a constant; $\tilde{\cal O}$ ignores polylogarithmic factors.
Method Sample complexity Step-size Sampling / Coverage
Simultaneous double $Q$-learning
Ours $\tilde{\mathcal{O}}\!\left(\frac{|\mathcal{S}\times\mathcal{A}|^{2}}{\varepsilon^{2}\,d_{\min}^{10}\,(1-\gamma)^{12}}\right)$ constant i.i.d., stochastic
Double $Q$-learning
Lin et al. [19] $\tilde{\mathcal{O}}\!\left(\frac{t_{\mathrm{cover}}}{(1-\gamma)^{7}\epsilon^{2}}\right)$ constant non-i.i.d., cover-time
Xiong et al. [18] $\tilde{\mathcal{O}}\!\left(\left(\frac{t_{\mathrm{cover}}^{4}}{(1-\gamma)^{6}\epsilon^{2}}\right)^{\!\frac{1}{\omega}}+\left(\frac{t_{\mathrm{cover}}^{2}}{1-\gamma}\right)^{\!\frac{1}{1-\omega}}\right)$ diminishing non-i.i.d., cover-time
Hasselt [14] – (asymptotic convergence only) diminishing i.i.d., infinite-time covering
Weng et al. [30] – (asymptotic convergence only) diminishing i.i.d., infinite-time covering
$Q$-learning
Lee et al. [13] $\tilde{\cal O}\!\left({\frac{{\gamma^{2}d_{\max}^{2}|{\cal S}\times{\cal A}|^{2}}}{{\varepsilon^{2}d_{\min}^{4}(1-\gamma)^{6}}}}\right)$ constant i.i.d., stochastic
Chen et al. [24] $\tilde{\cal O}\!\left(\frac{1}{d_{\min}^{3}(1-\gamma)^{5}\varepsilon^{2}}\right)$ constant non-i.i.d., stochastic
Li et al. [26] $\tilde{\cal O}\!\left(\frac{1}{d_{\min}(1-\gamma)^{5}\varepsilon^{2}}+\frac{t_{\rm mix}}{d_{\min}(1-\gamma)}\right)$ constant non-i.i.d., stochastic
Lim et al. [34] $\tilde{\mathcal{O}}\left(\frac{|\mathcal{S}\times\mathcal{A}|^{13}}{(1-\gamma)^{16}d_{\min}^{12}}\frac{1}{\epsilon^{2}}\right)$ diminishing non-i.i.d., stochastic
Beck et. al. [33] $\tilde{\cal O}\left({\frac{t_{\rm cover}^{3}|{\cal S}\times{\cal A}|}{(1-\gamma)^{5}\varepsilon^{2}}}\right)$ constant non-i.i.d., cover-time
Qu et. al. [25] $\tilde{\cal O}\left({\frac{{{t_{{\rm{mix}}}}|{\cal S}\times{\cal A}{|^{2}}}}{{{{(1-\gamma)}^{5}}{\varepsilon^{2}}}}}\right)$ diminishing non-i.i.d., cover-time
Even-Dar et. al. [8] $\tilde{\cal O}\left({\frac{{{{({t_{{\rm{cover}}}})}^{\frac{1}{{1-\gamma}}}}}}{{{{(1-\gamma)}^{4}}{\varepsilon^{2}}}}}\right)$ diminishing non-i.i.d., cover-time
Tsitsiklis [31] – (asymptotic convergence only) diminishing non-i.i.d., infinite-time covering
Jaakkola [32] – (asymptotic convergence only) diminishing non-i.i.d., infinite-time covering
Borkar et  al. [7] – (asymptotic convergence only) diminishing non-i.i.d., synchronous update
[…]
@@ 5 Framework for convergence analysis of SDQ
Before presenting the technical details, we briefly outline the main structure of the finite-time analysis. The central challenge in analyzing SDQ stems from the coupled and switching nature of the two estimators, which introduces additional affine terms and stochastic disturbances compared to standard $Q$-learning. To address this challenge, we first model SDQ as a discrete-time switching system. We then construct two auxiliary comparison systems—an upper comparison system and a lower comparison system—that respectively bound the original dynamics from above and below. The analysis proceeds by first controlling the evolution of the estimator disagreement $Q_{k}^{A}-Q_{k}^{B}$ through a dedicated error system. Once this disagreement is shown to contract over time, the lower comparison system effectively reduces to a stable linear stochastic system, enabling finite-time error bounds to be derived. Finally, combining the bounds from the comparison systems yields the finite-time expected error guarantees for SDQ. A detailed realization of this analysis plan, including the specific comparison systems and error dynamics, is provided in Section 6.1.
@@ 5.1 Switching system model
$w_{k}^{A}$ $=(e_{a_{k}}\otimes e_{s_{k}})r_{k}+\gamma(e_{a_{k}}\otimes e_{s_{k}})(e_{s_{k}^{\prime}})^{T}\Pi_{Q_{k}^{B}}Q_{k}^{A}-(e_{a_{k}}\otimes e_{s_{k}})(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}$
@@
$w_{k}^{B}$ $=(e_{a_{k}}\otimes e_{s_{k}})r_{k}+\gamma(e_{a_{k}}\otimes e_{s_{k}})(e_{s_{k}^{\prime}})^{T}\Pi_{Q_{k}^{A}}Q_{k}^{B}-(e_{a_{k}}\otimes e_{s_{k}})(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}$
@@
Here, $s_{k}\in\mathcal{S}$ and $a_{k}\in\mathcal{A}$ denote the state and action visited at iteration $k$, respectively, and $s_{k}^{\prime}\in\mathcal{S}$ denotes the subsequent state generated according to the transition probability $P(\cdot\mid s_{k},a_{k})$. Next, using the optimal Bellman equation $(\gamma DP\Pi_{Q^{*}}-D)Q^{*}+DR=0$ with (5.1), one can obtain
@@
which is a linear switching system with an extra affine terms, $\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q^{*}}Q^{*}\}$ and $\alpha D\{\gamma P\Pi_{Q_{k}^{A}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}\}$, and the stochastic noises, $w_{k}^{A}$ and $w_{k}^{B}$  [12]. The main difficulty in analysing the system arises from the extra affine term and the stochastic noise. Without these terms, finite-time analysis would be straightforward since the stability of the system matrix could be directly analyzed. However, with the affine term, the analysis becomes more challenging. To address this difficulty, we introduce the lower and upper comparison systems as in [12], which enable easier analysis.
@@ 5.2 Upper comparison system
$Q_{k+1}^{A_{U}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)(Q_{k}^{A_{U}}-Q^{*})+\alpha w_{k}^{A},\quad Q_{0}^{A_{U}}-Q^{*}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$,
$Q_{k+1}^{B_{U}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D)(Q_{k}^{B_{U}}-Q^{*})+\alpha w_{k}^{B},\quad Q_{0}^{B_{U}}-Q^{*}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$. (13)
Here, $Q_{k}^{A_{U}}$ and $Q_{k}^{B_{U}}$ denote the state–action value iterates of the upper comparison system associated with $Q_{k}^{A}$ and $Q_{k}^{B}$, respectively. The above systems are switching systems, which have system matrices $I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D$ and $I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D$. These matrices switch according to the changes of $Q_{k}^{A}$ and $Q_{k}^{B}$. We can prove that the trajectory of the upper comparison system bounds that of the original system from above.
@@ 5.3 Lower comparison system
$Q_{k+1}^{A_{L}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})+\alpha w_{k}^{A},\quad Q_{0}^{A_{L}}-Q^{*}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$,
$Q_{k+1}^{B_{L}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})+\alpha w_{k}^{B},\quad Q_{0}^{B_{L}}-Q^{*}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$, (14)
Here, $Q_{k}^{A_{L}}$ and $Q_{k}^{B_{L}}$ denote the state–action value iterates of the lower comparison system associated with $Q_{k}^{A}$ and $Q_{k}^{B}$, respectively. The stochastic noises $w_{k}^{A}$ and $w_{k}^{B}$ are identical to the original system (5.1). As before, we can prove that the trajectory of the lower comparison system bounds that of the original system from below.
@@
The proof is given in C.2. The lower comparison system (5.3) can be seen as a linear system with the states $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$ and the system matrix $I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D$. Moreover, it also includes the extra terms, $\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$ and $\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})$, which can be seen as external disturbances. To derive a finite-time error bound, one needs to establish bounds on the error $Q_{k}^{A}-Q_{k}^{B}$ first. Therefore, in the next subsections, we introduce an error system.
Figure 4: Overall flow of the proposed analysis
@@ 5.4 Error system
$Q_{k+1}^{\text{err}}$ $=(I-\alpha D)Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$, (15)
@@
However, the error system in (15) has the affine term $\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}$ similar to the original double $Q$-learning or $Q$-learning. Therefore, one can imagine that its convergence can be proved using similar techniques as in the $Q$-learning analysis [11, 12, 13]. In particular, one can derive the upper and lower comparison systems of the error system, where these two auxiliary systems respectively provide upper and lower bounds on the evolution of the error trajectory. This allows the overall convergence to be established by showing that the true error remains confined between the two comparison systems, and the lower comparison system is linear. However, for the error system (15), similar ideas cannot be applied because the lower comparison system of the error system is a switching system. For this reason, we will use a different approach.
To this end, we will first consider an upper comparison system of the error system, and then derive a lower comparison system of the upper comparison system, which is linear. Let $Q_{k}^{\text{err}_{U}}$ denote the state of the upper comparison system, and let $A_{Q_{k}^{\text{err}_{U}}}$ be its corresponding system matrix that depends on $Q_{k}^{\text{err}_{U}}$. Conceptually, one could analyze the convergence of upper comparison system by adapting a standard autocorrelation-based method, which tracks the evolution of the second moment $\mathbb{E}[Q_{k}^{\mathrm{err}_{U}}(Q_{k}^{\mathrm{err}_{U}})^{T}]$ using a linear recursion as shown in Lemma 3 of Appendix. However the present upper comparison system forms a switching system whose system matrix $A_{Q_{k}^{\mathrm{err}_{U}}}$ switches according to $Q_{k}^{\mathrm{err}_{U}}$ and depends probabilistically on its own state. Because of this dependence, taking expectation on both sides does not decouple the matrix and the state, and hence the simple linear recursion used in Lemma 3 cannot be directly applied to obtain $\mathbb{E}[Q_{k}^{\mathrm{err}_{U}}(Q_{k}^{\mathrm{err}_{U}})^{T}]$. To handle this coupling effect rigorously, we retain the switching-system-based analysis for $Q_{k}^{\mathrm{err}_{U}}$, which provides a mathematically clear characterization of its convergence behavior. Then, we will consider a lower comparison system of the error system. Because the lower comparison system is also a switching system, we will derive a subtraction system, which can be obtained by subtracting the error lower comparison system from the error upper comparison system.